OpenStudy (anonymous):

Solve the system of equations and choose the correct answer from the list of options. d + e = 12 –d + e = –8 Label the ordered pair as (d, e). A. (0, 0) B.(2, 10) C. (10, 2) D. (12, 0)

OpenStudy (mathstudent55):

Since you have d and -d, use the addition method.

OpenStudy (mathstudent55):

That will give the value of e easily. Then choose the only choice that has the correct value of e.

OpenStudy (anonymous):

d + e =12 -d + e = -8 ----------- 2e = 4 - - 2 2 ---------- e = 2 d + e = 12 d + 2 = 12 - 2 -2 ------------ d = 10

OpenStudy (anonymous):

So, (10,2)